# Stochastic Self-Organization in Multi-Agent Systems

### v1

@@ 2.2 Decentralized Initialization
This first stage of SelfOrg (referred to as collaboration round $t=0$) aims to generate a pool of diverse, but potentially noisy responses from $N$ agents. Given the user query $\mathcal{Q}$, each agent independently generates its own initial response $\mathcal{R}_{n}^{(0)}$. For this initial round, $\mathcal{P}_{n,\textrm{coll}}^{(0)}=\emptyset$ because agent $\mathcal{A}_{n}$ receives no input from other agents. We map each agent response $\mathcal{R}_{n}^{(0)}$ to an embedding $\mathbf{r}_{n}^{(0)}=f(\mathcal{R}_{n}^{(0)})$ with a lightweight model $f$ (e.g., all-MiniLM-L6 (Reimers & Gurevych, 2019)), which need not be the same LLM used by the agents. These embeddings provide a fixed-dimensional, semantically meaningful representation of the agent responses. Subsequent stages use these response embeddings to infer contributions and construct the communication graph.
@@ 2.3 Contribution Estimation
Let $L_{n}$ be the multiplicative factor from Theorem 1, and let $\underline{L}=\min_{m}L_{m}$. If
@@ 2.4 Communication Graph Formation
14:  Set $\mathcal{R}$$^{\star}$ as round output (fed to leading agent in next round)
@@
To avoid a fully connected graph, we retain only semantically meaningful links: for agent $\mathcal{A}_{n}$, an incoming candidate edge $e_{m\rightarrow n}^{(t+1)}\in\mathcal{E}^{(t+1)}$ is activated (set to $1$) if and only if $\mathbf{S}_{n,m}\geq\tau$, where $\tau$ is a similarity threshold and $\psi_{m}^{(t+1)}>\psi_{n}^{(t+1)}$. Alternatively, one may achieve sparsification by restricting active edges to $k$-most similar neighbors of each agent.
@@ 2.5 Response Propagation and Aggregation
Figure 2: Analysis of Qwen-1.5B agent over 100 runs on the same math problem (GSM-Hard).
@@ 2.6 Probabilistic Modeling of Multi-Agent System
Under Assumption 1, for every $n\in\mathcal{S}$ and $k\notin\mathcal{S}$ we have $\psi_{n}>\psi_{k}$, where $\psi_{n}=\cos(\mathbf{r}_{n},\mathbf{r}_{\mathrm{avg}})$ is the contribution score.
@@ 3.3 Heterogeneous Agents
We further evaluate configurations that mix strong and weak agents, with detailed results presented in Appendix D.1. Beyond accuracy, we also analyze efficiency in terms of token usage (Appendix D.2). Additional ablation studies examine the impact of the number of agents, the effect of reform across rounds, and the role of the embedding model in contribution estimation (Appendix E).
@@ 4 Related Work
Communication Graphs. Prior work has explored a spectrum of communication topologies. Fixed structures include chains, trees, complete graphs, and random graphs, with recent studies systematically comparing these patterns across task families such as mathematical reasoning, knowledge reasoning, and coding (Qian et al., 2025). Beyond static designs, some approaches treat the topology as optimizable: edges are sampled and trained with policy gradients or masks (Zhuge et al., 2024; Zhang et al., 2025a; Qian et al., 2025). A complementary line delegates topology design to a separate model that outputs a task- or query-specific communication graph (Zhang et al., 2025b; Ye et al., 2025b). Other frameworks rely on an external LLM “judge” to rank, filter, or finalize outputs (Liu et al., 2024; Zhang et al., 2025c; Zhuge et al., 2025; Ebrahimi et al., 2025). While effective in constrained settings, these strategies incur substantial overhead: pretraining graph generators, optimization over edges, or repeated calls to a judge LLM.
@@ A.1 Proof of Theorem 1
Let ${\bm{x}}=\sum_{m\in\mathcal{S}}\mathbf{r}_{m}$ and recall $\mathbf{r}_{\mathrm{avg}}=\frac{1}{N}\sum_{m=1}^{N}\mathbf{r}_{m}$.
@@ Appendix B Implementation Details
SelfOrg is configured as:
@@
  • Rounds and structure: maximum of $3$ rounds (including decentralized initialization); with DAG enforcement.
  • Contribution estimation: we use all-MiniLM-L6-v2 embedding model with embedding dimension of $384$ (lightweight sentence embedding model).
@@
  • Reform policy: we reform the DAG in each round from updated responses.
@@
We adopt a standard community template for defining agent roles, widely used in prior multi-agent system benchmarks. In our experiments, a subset of four/five agents is instantiated per run (default), selected in fixed order unless otherwise specified. Each role is assigned a default prompt template (system instruction) from the benchmark community, without additional fine-tuning or hand-engineering. This ensures that performance differences arise from orchestration rather than custom role design.
@@
We use a direct scoring approach using a task-specific evaluator (xVerify (Chen et al., 2025)), which is fine-tuned to assess correctness across various domains (Ye et al., 2025a).
@@ D.2 Token Consumption
D.3 Efficient SelfOrg
@@
For completeness, we provide Figures 9 and 10, which include efficient SelfOrg along with the other baseline methods and depict the performance and completion/prompt token consumption.
@@
D.4 Embedding Model
In our main experiments, we employ the all-MiniLM-L6-v2 (Reimers & Gurevych, 2019) model, a lightweight embedding model with only 22.722.7M parameters, to estimate similarity between agent responses. This choice is intentional: we aim to keep the method efficient and avoid reliance on large embedding models, even if this introduces some additional noise into similarity estimates.
To validate this design choice, we conduct an ablation study in the weak-agent-in-a-pool scenario using different embedding models. In addition to all-MiniLM, we evaluate all-MPNet-base-v2 (109109M parameters) (Reimers & Gurevych, 2019) and Qwen3-0.6B-Embedding (600600M parameters) (Zhang et al., 2025d). Across all cases, the embedding models are able to correctly identify the weakest agent: the weak participant is consistently ranked lowest in the majority of runs. Moreover, both MPNet and Qwen-0.6B provide sharper separation between strong and weak agents compared to MiniLM, reflecting their stronger representational capacity.
Appendix E Ablation Study
E.1 Number of Agents
@@
E.2 To Reform or Not To Reform
An important design choice in SelfOrg is whether to reform the communication graph between rounds of interaction. Reforming allows agents to dynamically update their information flow structure based on the latest responses, while a static graph keeps the initial topology fixed throughout. We conduct an ablation on two benchmarks, GSM8K and MMLU, using $N=5$ agents and neighbor budget $K=3$, to evaluate the impact of graph reform.
@@ E.2 To Reform or Not To Reform
As shown in Table 5, reforming the graph consistently improves performance, though the absolute gains are modest. This suggests that while the initial communication structure already captures useful alignment among agents, dynamically restructuring the graph allows the system to consolidate correct signals more effectively, especially on more challenging knowledge-intensive tasks. The relatively small gap also indicates that SelfOrg is robust to whether reform is applied, but benefits from it most in settings where agent responses are more diverse and noisy.

### v2

@@ 2.2 Decentralized Initialization
This first stage of SelfOrg (referred to as collaboration round $t=0$) aims to generate a pool of diverse, but potentially noisy responses from $N$ agents. Given the user query $\mathcal{Q}$, each agent independently generates its own initial response $\mathcal{R}_{n}^{(0)}$. For this initial round, $\mathcal{P}_{n,\textrm{coll}}^{(0)}=\emptyset$ because agent $\mathcal{A}_{n}$ receives no input from other agents. We map each agent response $\mathcal{R}_{n}^{(0)}$ to an embedding $\mathbf{r}_{n}^{(0)}=f(\mathcal{R}_{n}^{(0)})$ with a lightweight model $f$ (e.g., all-MiniLM-L6 (Reimers and Gurevych, 2019)), which need not be the same LLM used by the agents. These embeddings provide a fixed-dimensional, semantically meaningful representation of the agent responses. Subsequent stages use these response embeddings to infer contributions and construct the communication graph.
@@ 2.3 Contribution Estimation
Let $L_{n}$ be the multiplicative factor from Theorem 1, and let $\underline{L}=\min_{j}L_{j}$. If
@@ 2.4 Communication Graph Formation
14:  Set $\mathcal{R}$$^{\star}$ as round output (fed to the leading agent in next round)
@@
To avoid a fully connected graph, we retain only semantically meaningful links: for agent $\mathcal{A}_{n}$, an incoming edge $e_{m\rightarrow n}^{(t+1)}\in\mathcal{E}^{(t+1)}$ is activated if and only if $\mathbf{S}_{n,m}\geq\tau$, where $\tau$ is a similarity threshold and $\psi_{m}^{(t+1)}>\psi_{n}^{(t+1)}$. Alternatively, sparse connectivity may be achieved by limiting active edges to the $k$ most similar neighbors of each agent.
@@ 2.5 Response Propagation and Aggregation
Figure 2: Analysis of Qwen-1.5B over 100 runs on the same math problem (GSM-Hard).
@@ 2.6 Probabilistic Modeling of Multi-Agent System
Under Assumption 1, for every $n\in\mathcal{S}$ and $k\notin\mathcal{S}$ we have $\psi_{n}>\psi_{k}$, where $\psi_{n}=\cos(\mathbf{r}_{n},\mathbf{r}_{\mathrm{all}})$ is the contribution score.
@@ 3.3 Heterogeneous Agents
We further evaluate configurations that mix strong and weak agents, with detailed results presented in Appendix D.1. Beyond accuracy, we also analyze efficiency in terms of token usage (Appendix D.2). Additional ablation studies examine the impact of the number of agents, the effect of reform across rounds, and the role of the embedding model in contribution estimation (Appendix F).
@@ 4 Related Work
Communication Graphs. Prior work has explored a spectrum of communication topologies. Fixed structures include chains, trees, complete graphs, and random graphs, with recent studies systematically comparing these patterns across task families such as mathematical reasoning, knowledge reasoning, and coding (Qian et al., 2025). Beyond static designs, some approaches treat the topology as optimizable: edges are sampled and trained with policy gradients or masks (Zhuge et al., 2024; Zhang et al., 2025a; Qian et al., 2025). A complementary line delegates topology design to a separate model that outputs a task- or query-specific communication graph (Zhang et al., 2025b; Ye et al., 2025b). Other frameworks rely on an external LLM “judge” to rank, filter, or finalize outputs (Liu et al., 2024; Zhang et al., 2025c; Zhuge et al., 2025; Ebrahimi et al., 2025), or extend to decentralized settings (Yang et al., 2025; Lu et al., 2024). While effective in constrained settings, these strategies incur substantial overhead: pretraining graph generators, optimization over edges, or repeated calls to a judge LLM.
@@ 5 Conclusion
Acknowledgments
This material is partly based on work supported by the Office of Naval Research N00014-24-1-2168.
@@ A.1 Proof of Theorem 1
Let ${\bm{x}}=\sum_{j\in\mathcal{S}}\mathbf{r}_{m}$ and recall $\mathbf{r}_{\mathrm{avg}}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{r}_{m}$.
@@ Appendix B Implementation Details
Implementation can be found at: https://github.com/tnurbek/selforg. SelfOrg is configured as:
@@
  • Rounds and structure: a maximum of $3$ rounds (including decentralized initialization); with DAG enforcement.
  • Contribution estimation: we use all-MiniLM-L6-v2 embedding model with an embedding dimension of $384$ (a lightweight sentence embedding model).
@@
  • Reform policy: we reform the DAG in each round based on the updated responses.
@@
We adopt a standard community template for defining agent roles that is widely used in prior multi-agent system benchmarks. In our experiments, a subset of four/five agents is instantiated per run (default), selected in fixed order unless otherwise specified. Each role is assigned a default prompt template (system instruction) from the benchmark community, without additional fine-tuning or hand-engineering. This ensures that performance differences arise from orchestration rather than custom role design.
@@
We use a direct scoring approach with a task-specific evaluator (xVerify (Chen et al., 2025)), which is fine-tuned to assess correctness across various domains (Ye et al., 2025a).
@@ D.2 Token Consumption
Figure 8: Cosine-similarity heatmaps of response embeddings (correct vs. incorrect responses) across three benchmarks.
D.3 Probabilistic Modeling. Additional Correlation Results
To further substantiate the relationship between embedding similarity and answer correctness, we extend the empirical analysis beyond the illustrative instance (Figure 2) and evaluate the phenomenon across three benchmarks: GSM-Hard, AQUA-RAT, and MMLU in Figure 8. Across benchmarks, we observe a consistent geometric signature in the response-embedding space: embeddings of correct answers form a tight cluster, whereas incorrect answers are comparatively dispersed, with low cross-similarity between the two groups. This structure persists broadly across both open-ended and multiple choice settings.
Table 5: For randomly sampled queries, $p$ denotes the probability of generating the correct answer over $100$ runs; $\max_{k}p_{k}$ is the largest probability of incorrect answers.
GSM-Hard AQUA-RAT MMLU
$p$ $\max_{k}p_{k}$ $p$ $\max_{k}p_{k}$ $p$ $\max_{k}p_{k}$
11 $0.98$ $0.01$ $0.60$ $0.11$ $0.61$ $0.14$
22 $0.35$ $0.05$ $0.93$ $0.02$ $0.57$ $0.16$
33 $0.74$ $0.02$ $0.90$ $0.04$ $0.42$ $0.20$
44 $0.71$ $0.03$ $0.84$ $0.06$ $0.75$ $0.10$
55 $0.39$ $0.09$ $0.23$ $0.21$ $0.42$ $0.21$
66 $0.31$ $0.07$ $0.40$ $0.16$ $0.80$ $0.07$
77 $0.70$ $0.04$ $0.70$ $0.08$ $0.84$ $0.07$
In all three cases, (i) correct responses exhibit high within-group similarity, (ii) incorrect responses are more weakly correlated with each other, and (iii) cross-block similarities remain low. This pattern is consistent with the modeling assumption that correct responses are semantically aligned, while incorrect responses scatter across error modes, implying that correct responders tend to align more with the population centroid than incorrect responders.
To complement the qualitative heatmap plots, we randomly sample multiple queries from each task and estimate (over repeated stochastic generations) the probability of producing the correct answer, denoted by $p$, and the maximum probability of the most frequently occurring incorrect response, denoted by $\max_{k}p_{k}$. Table 5 reports these values. Across GSM-Hard, AQUA-RAT, and MMLU, $p$ consistently exceeds $\max_{k}p_{k}$ by a substantial margin (satisfying the assumption in Lemma 2), indicating that correctness concentrates probability mass more strongly than any competing incorrect alternative.
Notably, the narrowest separation appears in the 5th AQUA-RAT example (highlighted in red), where $p\approx 0.23$ and $\max_{k}p_{k}\approx 0.21$. This case is consistent with near-random behavior: AQUA-RAT has five answer options, so uninformed guessing concentrates around $p\approx 0.2$, and no semantically coherent correct cluster is expected to dominate in such instances. Outside of these near-chance cases, the observed gaps support the mechanism underlying the contribution dominance.
D.4 Embedding Model
In our main experiments, we employ the all-MiniLM-L6-v2 (Reimers and Gurevych, 2019) model, a lightweight embedding model with only 22.722.7M parameters, to estimate similarity between agent responses. This choice is intentional: we aim to keep the method efficient and avoid reliance on large embedding models, even if this introduces some additional noise into similarity estimates.
To validate this design choice, we conduct an ablation study in the weak-agent-in-a-pool scenario using different embedding models (see Figure 9). In addition to all-MiniLM, we evaluate all-MPNet-base-v2 (109109M parameters) (Reimers and Gurevych, 2019) and Qwen3-0.6B-Embedding (600600M parameters) (Zhang et al., 2025d). Across all cases, the embedding models are able to correctly identify the weakest agent: the weak participant is consistently ranked lowest in the majority of runs. Moreover, both MPNet and Qwen-0.6B provide sharper separation between strong and weak agents compared to MiniLM, reflecting their stronger representational capacity.
D.5 Heterogeneous Agents with Dynamic Roles
Figure 10: Dynamic role switching in SelfOrg with heterogeneous agents.
We additionally evaluate a dynamic role-switching variant of the heterogeneous-agent setting from Section 3.3 where agent roles are dynamic and not pre-fixed. Dynamic role settings yield a small improvement over fixed roles (accuracy: $66.93\%$ vs. $66.14\%$, Figure 4), suggesting that contribution-guided routing remains effective without a static role prior and can achieve similar performance with permutation diversity.
We present Figure 10, which reports the percentage of times each agents attains contribution ranks (141{-}4). The distribution indicates a systematic separation between stronger and weaker agents: Qwen appears most frequently in earlier positions (rank $1$), while Mistral is most frequently placed later (rank $4$).
Appendix E Efficient SelfOrg
@@
For completeness, we provide Figures 12 and 13, which include efficient SelfOrg along with the other baseline methods and depict the performance and completion/prompt token consumption.
@@
E.1 Termination Breakdown.
Efficient SelfOrg halts in one of two ways: (i) it reaches the pre-specified maximum number of communication rounds, or (ii) it stops early once semantic consensus is achieved, i.e., when the minimum pairwise cosine similarity among agent responses satisfies ${\mathbf{S}}_{\min}\geq\gamma$ (consensus threshold; as described above) before exhausting the round budget.
Table 6 reports the breakdown across datasets under two thresholds. With the looser consensus requirement $\gamma=0.9$, roughly $30{-}39\%$ of sessions terminate early; tightening the threshold to $\gamma=0.95$ reduces early terminations to about $11{-}15\%$. This behavior is consistent with the role of $\gamma$ as an efficiency parameter: stricter value of $\gamma$ demands stronger agreement (fewer early stops), while looser values trigger early stopping more often, reducing token usage while keeping accuracy comparable in aggregate.
Table 6: Fraction of runs terminating by round limit (“All rounds”) versus early stopping due to consensus ($\gamma$) under two thresholds.
$\gamma$ $\gamma=\mathbf{0.9}$ $\gamma=\mathbf{0.95}$
Dataset All rounds Early stopped All rounds Early stopped
AQUA-RAT $0.669$ $0.331$ $0.846$ $0.154$
MATH $0.612$ $0.388$ $0.852$ $0.148$
MMLU $0.700$ $0.300$ $0.894$ $0.106$
Appendix F Ablation Study
F.1 Number of Agents
@@
F.2 To Reform or Not To Reform
An important design choice in SelfOrg is whether to reform the communication graph between rounds of interaction. Reforming allows agents to dynamically update their information flow structure based on the latest responses, while a static graph keeps the initial topology fixed throughout. We conduct an ablation study on two benchmarks, GSM8K and MMLU, using $N=5$ agents and a neighbor budget of $K=3$, to evaluate the impact of graph reform.
@@ F.2 To Reform or Not To Reform
As shown in Table 7, reforming the graph consistently improves performance, though the absolute gains are modest. This suggests that while the initial communication structure already captures useful alignment among agents, dynamically restructuring the graph allows the system to consolidate correct signals more effectively, especially on more challenging knowledge-intensive tasks. The relatively small gap also indicates that SelfOrg is robust to whether reform is applied but benefits from it most in settings where agent responses are more diverse and noisy.
F.3 Aggregation Strategies
We evaluate the sensitivity of SelfOrg to the final-answer aggregation rule by substituting several alternative methods while keeping the rest of the method unchanged. Table 8 reports the performance on MMLU over 3 runs. The contribution-weighted centroid used in SelfOrg is competitive with selecting the highest contributor across all rounds, and both outperform the uniform centroid and voting-based baselines. Naive majority voting suffers a large drop and high variance due to option-letter parsing errors; using an LLM to adjudicate the majority recovers most of this gap but introduces additional variance.
Table 8: Ablation study on the aggregation methods on MMLU (mean $\pm$ std over 33 runs). Naive majority voting over extracted option letters; parsing errors materially degrade performance and increase variance.
Aggregation method MMLU [mean $\pm$ std]
Contribution-weighted centroid (SelfOrg , default) $\mathbf{53.60\pm 1.27}$
Uniform centroid $52.70\pm 1.31$
Majority voting $46.17\pm 5.35$
Majority voting (with LLM deciding) $51.90\pm 2.33$
Highest contributor across all rounds $53.20\pm 1.19$
F.4 SelfOrg with Memory and Tools
All main-table results are reported in a tool-agnostic setting to isolate orchestration quality. Here we show that SelfOrg’s response-conditioned routing remains effective when agents are augmented with (i) per-agent memory (a scratchpad persisted across rounds) and (ii) a lightweight tool (sandboxed Python REPL for arithmetic checks; no external access). The orchestration mechanism is unchanged: contributions are still computed from the semantic content of agent responses and used to form the response-conditioned DAG.
Table 9: Effect of memory and tools on SelfOrg with Qwen-1.5B backbone.
Method (Qwen-1.5B) GSM-Hard MATH
Single 36.2 49.2
Self-Consistency 37.2 50.2
SelfOrg (base; no memory/tools) 38.0 52.4
SelfOrg + memory 37.8 53.4
SelfOrg + memory + tools 40.0 54.2
Table 9 compares single call, self-consistency (Wang et al., 2023), base SelfOrg , and SelfOrg with memory/tools under Qwen-1.5B. Base SelfOrg improves upon both single and self-consistency methods. Adding memory yields comparable performance on GSM-Hard and a small gain on MATH, while adding both memory and tools provides the largest gains on both tasks.
Appendix G Additional Coding Experiments
To evaluate whether SelfOrg extends beyond AQ/reasoning benchmarks, we add two coding evaluations using a weak backbone (Qwen-1.5B), where orchestration effects are expected to be most pronounced. We report HumanEval (Chen et al., 2021) with execution-based pass@1@1 and (ii) SRDD (Qian et al., 2024b) using the benchmark’s composite score (completeness, executability, consistency) since SRDD has no single ground-truth solution.
Table 10: Collaborative coding results with Qwen-1.5B backbone model (mean $\pm$ std). HumanEval is pass@1@1; SRDD is the benchmark composite score.
Method HumanEval SRDD
Single $47.31\pm 0.91$ $55.40\pm 0.72$
CoT $46.34\pm 0.75$ $55.07\pm 0.76$
Self-Consistency $45.63\pm 1.41$ $53.33\pm 0.76$
DyLAN $42.37\pm 3.47$ $37.40\pm 0.80$
MacNet $44.67\pm 0.80$ $56.00\pm 0.40$
SelfOrg $\mathbf{48.97\pm 0.89}$ $\mathbf{60.13\pm 0.90}$
SelfOrg is best on both HumanEval and SRDD, indicating that its contribution-guided communication improves not only reasoning but also collaborative coding.